# Gravity from equilibrium thermodynamics of stretched light cones

### v2

@@ III.2 $\theta$, $\sigma_{\mu\nu}$, and $\omega_{\mu\nu}$ for SLCs
We split the velocity into its curved and flat contribution as follows
@@
where $u^{\mu}_{f}$ is the flat contribution already defined in equation (7), and $u^{\mu}_{c}$ is an unknown curved contribution that vanishes in flat spacetime. We need to define $u^{\mu}_{c}$ in a way that keeps the complete velocity field in equation (15), normalized to $g_{\mu\nu}u^{\mu}u^{\nu}=-1$. Since we are only interested in studying the congruence generated by $u^{\mu}$ for small values of $t$, we can conveniently expand $u^{\mu}$ in powers of $t/\alpha\leq\epsilon/\alpha\ll 1$, keeping up to second order terms, i.e.:
@@ V Entropy balance
that are the sought for Einstein’s equations [^2] On the same grounds as in the previous work of three of the authors [11], it can be argued that the resulting gravitational dynamics are consistent with Weyl transverse gravity rather than with general relativity. Since this difference is unimportant for the purposes of the present paper, we do not discuss it here..

### v4

@@ III.2 $\theta$, $\sigma_{\mu\nu}$, and $\omega_{\mu\nu}$ for SLCs
To proceed, we first need to understand the curved spacetime construction of velocity $u^{\mu}$. Using the Riemann normal coordinate system, we can easily introduce the flat spacetime velocity $u^{\mu}_{f}$ given by equation (7) in any regular point $p$ of a generic curved spacetime [^1] The introduction of $u^{\mu}_{f}$ in a curved spacetime does not depend on using the Riemann normal coordinates in any way, they just make its construction particularly convenient.. However, due to curvature-dependent corrections to the metric (captured by equation (11)) $u^{\mu}_{f}$ only satisfy the normalization condition $g_{\mu\nu}u^{\mu}_{f}u^{\nu}_{f}=-1$, directly in $p$, and not in its vicinity. The total velocity $u^{\mu}$ must then also include curvature-dependent corrections. Therefore, we define it as follows
@@
where $u^{\mu}_{f}$ is the flat contribution, and $u^{\mu}_{c}$ is an unknown curved contribution that vanishes in flat spacetime. We need to define $u^{\mu}_{c}$ in a way that keeps the complete velocity field in equation (15), normalized to $g_{\mu\nu}u^{\mu}u^{\nu}=-1$. Since we are only interested in studying the congruence generated by $u^{\mu}$ for small values of $t$, we can conveniently expand $u^{\mu}$ in powers of $t/\alpha\leq\epsilon/\alpha\ll 1$, keeping up to second order terms, i.e.:
@@ V Entropy balance
Figure 2: Flow diagram of the derivation of Einstein equations from the Clausius equilibrium relation. Blue boxes in the diagram correspond to matter contributions, yellow boxes to geometry contributions, and green boxes to their combination.
@@
that are the sought for Einstein’s equations [^3] On the same grounds as in the previous work of three of the authors [11], it can be argued that the resulting gravitational dynamics are consistent with Weyl transverse gravity rather than with general relativity. Since this difference is unimportant for the purposes of the present paper, we do not discuss it here.. For convenience, we summarize the basic logic of our derivation of equations (73) from equilibrium thermodynamics in the flow diagram 2.
@@ VI Conclusions
Our work focused on the leading order contributions to the Clausius equilibrium relation in powers of the SLC’s size parameter $\alpha$ and the short time scale $\epsilon$. Since we consider both $\alpha$ and $\epsilon$ to be much smaller than the local curvature length scale, they certainly represent only small corrections to the leading order terms we studied. Nevertheless, a natural question is whether these subleading contributions can encode corrections to the Einstein’s equations. Two of us have previously studied this question in the vacuum setting [36]. We found that, since these correction terms include four or more derivatives of the metric, for dimensional reasons they would appear in the gravitational equations multiplied by suitable powers of $\alpha$ and/or $\epsilon$. However, as we discussed, both of these length scales are arbitrary (within a given range) in our treatment. Hence, the gravitational dynamics cannot depend on them. To study the corrections to the Einstein’s equations in this way, one would need to introduce a physically motivated length scale (e.g. in the way we explored here [36]).